OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. D(5, -2) m = 2/5

OpenStudy (anonymous):

For the equation y = mx+c, m = slope. Now, you have the slope, so it becomes y= (2/5)x +c -(1) Put (5, -2) into (1) -2 = 2/5 ( 5) + c -2 = 2 + c c = -4 Can you rewrite the equation now?

OpenStudy (anonymous):

i dont know how..

OpenStudy (anonymous):

y = mx+c What is m? what is c?

OpenStudy (anonymous):

m=2/5 c=-4

OpenStudy (anonymous):

i need to put it in this form.. _x_ _ y_____

OpenStudy (anonymous):

y = mx+c y = 2/5 (x) -4 Now, rearrange the terms... Multiply both sides by 5, what do you get?

OpenStudy (anonymous):

5y=2-20

OpenStudy (anonymous):

.

OpenStudy (anonymous):

Oh.. You missed the x.. 5y=2x-20 Now, subtract 5y from both sides, what do you get?

OpenStudy (anonymous):

but,...

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

i need to put it in this form.. _x_ _ y_____

OpenStudy (anonymous):

-5y?

OpenStudy (anonymous):

Do you need it in the form _x __y = _ or _x __y ___ = 0 ?

OpenStudy (anonymous):

Yes, -5y from both sides

OpenStudy (anonymous):

any form

OpenStudy (anonymous):

Okay... -5y from both sides, what do you get?

OpenStudy (anonymous):

5y=2x-20 5y - 5y = 2x - 20 - 5y

OpenStudy (anonymous):

2x-5y=-20?

OpenStudy (anonymous):

Nope... 5y - 5y = 2x - 20 - 5y 0 = 2x - 5y - 20 Got it?

OpenStudy (anonymous):

yup:D thanks

OpenStudy (anonymous):

Welcome :)